# Gravity and fluid dynamic correspondence on a null hypersurface: inconsistencies and advancement

### v1

@@ 3.2 BY-like tensor on a null surface
The above formulation of BY tensor and thereby defining the quasilocal BY parameters has been previously formulated in [51, 52], which is valid for the timelike surfaces. Since our focus is on the null surface, we look for the analogous term of the BY tensor for the null surface. However, as discussed earlier, obtaining a BY-like tensor in a direct manner (i.e. conjugate quantity of the induced metric) is not possible for the null surface as one cannot define an induced metric in this case. In addition, the suitable boundary term for the null surface was not known until the recent works [39, 40, 57, 58, 59, 60, 61]. In the following, we define a BY-like tensor in the following, which has been discussed in the recent papers [41, 42, 53].
@@
$\delta\mathcal{A}_{WP}^{(null)}=\delta\mathcal{A}_{EH}+\delta\mathcal{A}_{boundary}^{(null)}=\frac{1}{16\pi}\delta\int\sqrt{-g}R~{}d^{4}x+\frac{1}{8\pi}\delta\int_{\mathcal{H}}\sqrt{-g}\Big{(}\theta+\kappa\Big{)}d^{3}x~{},$
@@
Interestingly if one investigates the projection $q^{a}_{c}G_{ab}l^{b}$, which upon use of Einstein’s equation leads to famous DNS equation 27, includes the information of above $\mathcal{P}_{ab}$ and $\mathcal{P}^{(l)}_{a}$. Indeed the DNS equation can be expressed in terms of these conjugate momenta. Following the details of Appendix A, one can show that
@@
Use of these in 27 one finds a new form of DNS equation as
@@
$T^{a}_{~{}~{}b}|_{(null)}=2q^{ai}\mathcal{P}_{ib}+l^{a}\mathcal{P}^{(l)}_{b}=\frac{1}{8\pi}\Big{(}W^{a}_{~{}~{}b}-\Pi^{a}_{~{}b}W\Big{)}~{},$ (39)
where $W^{a}_{~{}~{}b}=\theta^{a}_{b}+l^{a}\omega_{b}$ and $W=\theta+\kappa$ The above expression of BY-like tensor in 39 resembles to the BY tensor of the timelike surface where $K^{a}_{b}$ is replaced by $W^{a}_{~{}~{}b}$ and $h^{a}_{b}$ is replaced by $\Pi^{a}_{~{}b}$ In the following, we show that this is the same quantity, which we are looking after.
@@ 4 Hydrodynamic description on a null surface: A formulation based on EM tensor
Let us now establish a one-to-one correspondence between the fluid energy-momentum tensor and the null Brown-York tensor which has been defined in the previous section. At first, taking a cue from fluid dynamics (see the text below 1), we identify the energy, momentum and spatial stress from different components of the null BY-like tensor, which can be obtained as
$\epsilon_{\rm null}=-T^{a}_{~{}~{}b}|_{(null)}k_{a}l^{b}=-\frac{\theta}{8\pi}~{},$
$p^{c}=-T^{a}_{~{}~{}b}|_{(null)}k_{a}q^{bc}=-\frac{\Omega^{c}}{8\pi}~{},$
$s^{ab}=q^{a}_{c}q^{bd}T^{c}_{~{}~{}d}|_{(null)}=2\eta\sigma^{ab}+q^{ab}(\zeta\Theta-P)~{},$ (40)
@@
where $\Omega^{a}=q^{ab}\omega_{b}$, $\eta=1/16\pi$, $\zeta=-1/16\pi$, $\sigma^{ab}=\theta^{ab}-q^{ab}\theta/2$, $\Theta=\theta$ and $P=\kappa/8\pi$. In what follows, we shall check the consistency of such identifications. Firstly, note that one can show (for details, see B)
$\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=\frac{1}{8\pi}R_{ab}\Pi^{a}_{~{}c}l^{b}~{},$ (41)
@@
where the differentiation operator $\mathcal{D}_{a}$ is compatible with the projection tensor $\Pi^{a}_{~{}b}$ i.e. $\mathcal{D}_{a}\Pi^{a}_{~{}b}=0$ where $\mathcal{D}_{a}A^{b}=\Pi^{i}_{~{}a}\Pi^{b}_{~{}j}\nabla_{i}(\Pi^{j}_{~{}k}A^{k})$ and $\mathcal{D}_{a}B_{b}=\Pi^{i}_{~{}a}\Pi^{j}_{~{}b}\nabla_{i}(\Pi^{k}_{~{}j}B_{k})$. Thus, in the absence of the external matter field, upon use of Einstein’s equation one obtains $\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=0$, which is analogous to the conservation of fluid energy-momentum tensor as described in 1. Note that such is conserved upon use of Einstein’s equation of motion and therefore the conservation of EM tensor carries the dynamics of gravity. In addition, from the above relation 41, we can obtain the following results for the gravitational fluid dynamics. Firstly, contracting 41 with the auxiliary null vector one finds
$k^{c}\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=0~{}.$ (42)
@@
The above equation indicates that the projection of 41 along the auxiliary null vector vanishes even without using the dynamical information of the null surface (i.e. without using the equation of motion). Hence, it can be considered as a mere geometrical identity of a null surface. Since it does not carry any dynamical information, we do not expect to have any significance from the viewpoint of null hydrodynamics, as we expect the latter to be related to the dynamics of the null hypersurface. On the contrary, contracting 41 with the induced metric of $\mathcal{B}^{(null)}$ provides
$q^{c}_{d}\mathcal{D}_{a}T^{a}_{~{}~{}c}|_{(null)}=\frac{1}{8\pi}R_{ab}q^{a}_{d}l^{b}~{}.$ (43)
Use of vacuum Einstein’s equation leads to vanishing of the right hand side and then it must be analogous to 3. Therefore we expect that the above one must give dynamics of fluid-description of gravity. Upon expanding the left-hand side of 43 and replacing $R_{ab}$ with the energy-momentum tensor corresponding to the external matter field (i.e.$G_{ab}=8\pi GT_{ab}^{(ext)}$), we obtain the DNS equation 27. The important point to note is that we have obtained the DNS equation (i.e. 27) conventionally by taking the appropriate projection of the conservation of energy-momentum tensor. In other words, 27 is analogous to 3. In addition, the fluid parameters are also identified in a conventional manner, i.e. from the components of the null BY tensor. For a long time, 27 has been correlated with the momentum-conservation relation of the Navier-Stokes equation, i.e. 8, where we find that our earlier identifications in 40 become consistent. With our previous identification, in the absence of external matter, 27 looks as
$q^{n}_{a}\pounds_{l}p_{n}-\Theta p_{n}-\zeta~{}^{(2)}D_{a}\Theta+~{}^{(2)}D_{a}P-2\eta~{}^{(2)}D_{i}\sigma^{i}_{a}=0~{},$ (44)
where, $P$ can be identified as $P=\kappa/8\pi$. Also, the shear viscosity coefficient ($\eta$) and the bulk viscosity coefficient ($\zeta$) can be identified as $\eta=1/16\pi$ and $\zeta=-1/16\pi$. The above equation, famously known as the Damour-Navier-Stokes equation, resembles to the momentum conservation equation of fluid dynamics (i.e. 8). The only difference, however, is the presence of the Lie-derivative of the momentum density (instead of the convective derivative of 8). Upon using Padmanabhan’s prescription of local inertial frame in adapted coordinates (see the metric 16) and a constraint $\partial_{t}q_{AB}=0$, 27 becomes (in the absence of external matter)

### v2

@@ 3.2 BY-like tensor on a null surface
The above formulation of BY tensor and thereby defining the quasilocal BY parameters has been previously formulated in [56, 57], which is valid for the timelike surfaces. Since our focus is on the null surface, we look for the analogous term of the BY tensor for the null surface. However, as discussed earlier, obtaining a BY-like tensor in a direct manner (i.e. conjugate quantity of the induced metric) is not possible for the null surface as one cannot define an induced metric in this case. In addition, the suitable boundary term for the null surface was not known until the recent works [62, 63, 64, 65, 66, 67]. In the following, we define a BY-like tensor in the following, which has been discussed in the recent papers [68, 53, 58].
@@
$\delta\mathcal{A}_{\rm WP}^{\rm(null)}=\delta\mathcal{A}_{\rm EH}+\delta\mathcal{A}_{\rm boundary}^{\rm(null)}=\frac{1}{16\pi}\delta\int\sqrt{-g}R~d^{4}x+\frac{1}{8\pi}\delta\int_{\mathcal{H}}\sqrt{-g}\Big(\theta+\kappa\Big)d^{3}x~,$
@@
Interestingly if one investigates the projection $q^{a}_{c}G_{ab}l^{b}$, which upon use of Einstein’s equation leads to famous DNS equation Eq. (, includes the information of above $\mathcal{P}_{ab}$ and $\mathcal{P}^{(l)}_{a}$. Indeed the DNS equation can be expressed in terms of these conjugate momenta. Following the details of Appendix Section A, one can show that
@@
Use of these in Eq. ( one finds a new form of DNS equation as
@@
$T^{a}_{~~b}|_{\rm(null)}=2q^{ai}\mathcal{P}_{ib}+l^{a}\mathcal{P}^{(l)}_{b}=\frac{1}{8\pi}\Big(W^{a}_{~~b}-\Pi^{a}_{~b}W\Big)~,$ (31)
where $W^{a}_{~~b}=\theta^{a}_{b}+l^{a}\omega_{b}$ and $W=\theta+\kappa$ . The above expression of BY-like tensor in Eq. ( resembles to the BY tensor of the timelike surface where $K^{a}_{b}$ is replaced by $W^{a}_{~~b}$ and $h^{a}_{b}$ is replaced by $\Pi^{a}_{~b}$ . In the following, we show that this is the same quantity, which we are looking after.
@@ 4 Hydrodynamic description on a null surface: A formulation based on EM tensor
Let us now establish a one-to-one correspondence between the fluid energy-momentum tensor and the null Brown-York tensor which has been defined in the previous section. At first, taking a cue from fluid dynamics (see the text below Eq. (), we identify the energy, momentum and spatial stress from different components of the null BY-like tensor, which can be obtained as
$\epsilon_{\rm null}=-T^{a}_{~~b}|_{\rm(null)}k_{a}l^{b}=-\frac{\theta}{8\pi}~,$
$p^{c}=-T^{a}_{~~b}|_{\rm(null)}k_{a}q^{bc}=-\frac{\Omega^{c}}{8\pi}~,$
$s^{ab}=q^{a}_{c}q^{bd}T^{c}_{~~d}|_{\rm(null)}=2\eta\sigma^{ab}+q^{ab}(\zeta\theta-P)~,$ (32)
@@
where $\Omega^{a}=q^{ab}\omega_{b}$, $\eta=1/16\pi$, $\zeta=-1/16\pi$, $\sigma^{ab}=\theta^{ab}-q^{ab}\theta/2$, $\theta=\theta$ and $P=\kappa/8\pi$. In what follows, we shall check the consistency of such identifications. Firstly, note that one can show (for details, see Section B) [^5] Here $\mathcal{D}^{a}$ represents the differentiation operator that is compatible with the projection tensor $\Pi^{a}_{~b}$. $\mathcal{D}_{a}A^{b}=\Pi^{i}_{~a}\Pi^{b}_{~j}\nabla_{i}(\Pi^{j}_{~k}A^{k})$ and $\mathcal{D}_{a}B_{b}=\Pi^{i}_{~a}\Pi^{j}_{~b}\nabla_{i}(\Pi^{k}_{~j}B_{k})$. Hence, it can be proved that $\mathcal{D}_{a}\Pi^{a}_{~b}=0$, implying that $\mathcal{D}_{a}$ is compatible with the projection tensor $\Pi^{a}_{~b}$
$\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=\frac{1}{8\pi}R_{ab}\Pi^{a}_{~c}l^{b}~,$ (33)
@@
Thus, in the absence of the external matter field, upon use of Einstein’s equation one obtains $\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=0$, which is analogous to the conservation of fluid energy-momentum tensor as described in Eq. (. Note that such is conserved upon use of Einstein’s equation of motion and therefore the conservation of EM tensor carries the dynamics of gravity. In addition, from the above relation Eq. (, we can obtain the following results for the gravitational fluid dynamics. Firstly, contracting Eq. ( with the auxiliary null vector one finds
$k^{c}\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=0~.$ (34)
@@
The above equation indicates that the projection of Eq. ( along the auxiliary null vector vanishes even without using the dynamical information of the null surface (i.e. without using the equation of motion). Hence, it can be considered as a mere geometrical identity of a null surface. Since it does not carry any dynamical information, we do not expect to have any significance from the viewpoint of null hydrodynamics, as we expect the latter to be related to the dynamics of the null hypersurface. On the contrary, contracting Eq. ( with the induced metric of $\mathcal{B}^{\rm(null)}$ provides
$q^{c}_{d}\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=\frac{1}{8\pi}R_{ab}q^{a}_{d}l^{b}~.$ (35)
[…]
$q^{\rm N}_{a}\pounds_{l}p_{n}-\theta p_{n}-\zeta~^{(2)}D_{a}\theta+~^{(2)}D_{a}P-2\eta~^{(2)}D_{i}\sigma^{i}_{a}=0~,$ (36)
where, $P$ can be identified as $P=\kappa/8\pi$. Also, the shear viscosity coefficient ($\eta$) and the bulk viscosity coefficient ($\zeta$) can be identified as $\eta=1/16\pi$ and $\zeta=-1/16\pi$. The above equation, famously known as the Damour-Navier-Stokes equation, resembles to the momentum conservation equation of fluid dynamics [25]. The only difference, however, is the presence of the Lie-derivative of the momentum density instead of the convective derivative. In the following section, when the above equation Eq. ( is re-written in terms of the adapted coordinate, it can be shown that the same equation (with some conditions) provides a direct resemblance with the momentum conservation equation of a non-relativistic fluid. Using the identifications of the fluid parameters, given in Eq. (, one can find the exact values of the same for the null-null structure of a spacetime. To illustrate this in Section C we calculate the fluid parameters corresponding a null surface defined on Schwarzschild, Kerr and FLRW spacetimes, respectively.
Now the question arises whether there exists an analogous relation to the energy–conservation equation for null hydrodynamics. To investigate this, we take another projection of Eq. (, contracting it with the null generator $l^{c}$, which gives
$l^{c}\mathcal{D}_{a}T^{a}_{~~c}|_{\rm(null)}=\frac{1}{8\pi}R_{ab}l^{a}l^{b}~.$ (37)
Expanding the left-hand side yields the Raychaudhuri equation Eq. (, a classical geometrical identity in general relativity. This equation has long been employed in the proof of the black-hole area theorem and hence occupies a central role in the thermodynamic interpretation of horizons. Here, however, we show that it also admits a fluid interpretation. In vacuum the right-hand side of (Eq. () vanishes, so the equation becomes directly analogous to (Eq. (). Therefore we expect it to be interpretable as an energy–conservation relation within the fluid description of gravity. This correspondence is particularly transparent in adapted coordinates; in the next section we will demonstrate that Eq. ( indeed takes the form of a fluid energy–conservation equation. Thus, in a fully covariant manner, we connect the gravitational dynamics on $\mathcal{H}$ with fluid dynamics beginning from a fluid energy–momentum tensor in the null-foliation setting.
Before proceeding, we remark on the distinction between our approach and the recent literature on Carrollian fluids. It should be pointed out that the previous works, which interprets in terms of Carroll fluid, have been done under a different geometrical construction. That particular construction is essentially that of a stretched horizon, employed both in the membrane and the Carroll paradigms, where the spacetime about the internal null boundary $\mathcal{H}$ is foliated by a stack of timelike surfaces. Now, under this stretched horizon purview, the analysis leads to a fluid description of the null surface [69, 55, 70, 39]. As can be evident, these constructions rely on the existence of a family of timelike stretched surfaces/horizons, with the null hypersurface being a limiting case in this family of stretched horizons. That is, geometrically the neighborhood of the null surface is foliated by a family of timelike hypersurfaces (as opposed to null surfaces as in our case).
In the Carrollian membrane paradigm, the vicinity of the null boundary $\mathcal{H}$ (located at a finite distance) is foliated by a family of three-dimensional timelike surfaces $\mathcal{T}$. Such a family of timelike surfaces is defined by the scalar function $r(x^{a})=\text{constant}>0$. The null boundary is obtained in the limit $r(x^{a})=0$. That is, essentially, the null limit to $\mathcal{H}$ is obtained by taking the limit $r\rightarrow 0$. However, such a family $\mathcal{T}$ is not foliated by spacelike hypersurfaces $\Sigma_{t}$ (which would in turn provide the notion of a time evolution vector field $\bm{t}$ and hence predispose the structures $\mathcal{H}$ and $\mathcal{T}$ with a Galilean picture). Instead, conformal Carroll structures are imposed on these timelike surfaces $\mathcal{T}$ as well as the null boundary $\mathcal{H}$ [39, 43]. The normal form to these Carroll horizons $\mathcal{T}$ is given by
$\underline{\bm{s}}=e^{\bar{\alpha}}\underline{\bm{d}}r~,$ (38)
with $\tilde{\rho}$ being a smooth scalar in $(\mathcal{M},\textbf{g},\bm{\nabla})$. The Carroll structure becomes null only on the null boundary obtained by taking $\tilde{\rho}=0$. In that case, the normal vector $\bm{s}$ coincides with the null generators $\bm{l}$ of $\mathcal{H}$. With such a Carrollian structure implemented on $\mathcal{T}$, an appropriate Carrollian fluid energy-momentum tensor can be described on the stretched horizon [39, 43].
At this point, it becomes quite imperative to discuss the differences in the approach of stretched membranes and those employed in our present study. At the very outset, the obvious constructional difference is regarding the foliation of $(\mathcal{M},\textbf{g},\bm{\nabla})$ about the null boundary $\mathcal{H}$. The stretched surface paradigm involves foliating the spacetime in the vicinity of $\mathcal{H}$ by timelike hypersurfaces. In our work, following Carter, we impose a null foliation of spacetime. By this, we avoid the issues of singularity that arise in the membrane paradigm picture by taking the null limit. However, as we show in the next section, in line with the membrane paradigm [69, 55, 70], we can employ a $3+1$ decomposition of the null boundary $\mathcal{H}$. This gives rise to a preferred time evolution vector field $\bm{t}$ take breaks general covariance bringing into the structure of $\mathcal{H}$, the Galilean picture. This is avoided in the Carrollian paradigm. Another minor technical difference that arises in the constructions is that of the nature of the null rigging vector $\bm{k}$. For stretched horizon method, the null rigging vector is assumed to generate null geodesics i.e. they satisfy the auto-parallel equation ($\nabla_{\bm{k}}\bm{k}=\tilde{\kappa}\bm{k}$, $\tilde{\kappa}$ being the non-affinity parameter). This is also the same case as employed in the construction of Gaussian null coordinates w.r.t. the null boundary $\mathcal{H}$. However, in our case of the null foliation of spacetime, the auxiliary null vector field is not constrained to be geodesic.
5 Is there any non-relativistic counterpart?
$D^{\rm N}\rho+\rho\theta^{\rm N}=0~,$
$\rho D^{\rm N}\epsilon^{\rm N}+P\theta^{\rm N}+\partial_{\mu}Q^{\mu}-\zeta(\theta^{\rm N})^{2}-2\eta(\sigma^{\rm N})^{\mu\nu}\sigma^{\rm N}_{\mu\nu}=0~,$
$\partial^{\mu}P+\rho D^{\rm N}v^{\mu}-\partial_{\nu}\Big[\zeta\theta^{\rm N}\delta^{\mu\nu}+2\eta(\sigma^{\rm N})^{\mu\nu}\Big]=0~,$ (40)
where
$D^{\rm N}\equiv\partial_{t}+v^{\mu}\partial_{\mu}~,$
$\sigma^{\rm N}_{\mu\nu}\equiv\partial_{(\mu}v_{\nu)}-\frac{1}{3}\delta_{\mu\nu}\theta^{\rm N}~,$
$\theta^{\rm N}=\partial_{\mu}v^{\mu}~.$ (41)
Here the index “N” is used for Navier-Stokes (non-relativistic) fluid, which is used to distinguish the physical quantities from the gravitational counterpart. Furthermore, $\rho$ is the mass density of the fluid (mass per unit volume), $v^{\mu}$ represents the 3-velocity components of the fluid ($\mu=1,2,3$), $P$ is the thermodynamic pressure, $\epsilon^{\rm N}$ represents specific internal energy (per unit mass) in the non-relativistic limit, $Q^{\mu}$ is the heat flux vector, $\zeta$ is the bulk viscosity coefficient, $\eta$ is the shear viscosity coefficient, $\theta^{\rm N}$ represents the expansion scalar, $\sigma^{\rm N}_{\mu\nu}$ represents the shear viscosity tensor, and $D^{\rm N}$ corresponds the convective derivative. If one uses the continuity equation in the energy conservation and momentum conservation relation, one obtains the modified form of energy and momentum conservation relations, where the energy conservation relation is given as
$D^{\rm N}(\rho\epsilon^{\rm N})+\Big(\rho\epsilon^{\rm N}+P\Big)\theta^{\rm N}+\partial_{\mu}Q^{\mu}-\zeta(\theta^{\rm N})^{2}-2\eta(\sigma^{\rm N})^{\mu\nu}\sigma^{\rm N}_{\mu\nu}=0~,$ (42)
and the momentum conservation relation is given as
$D^{\rm N}\Big(\rho v^{\mu}\Big)+\rho v^{\mu}\theta^{\rm N}+\partial^{\mu}P-\partial_{\nu}\Big[\zeta\theta^{\rm N}\delta^{\mu\nu}+2\eta(\sigma^{\rm N})^{\mu\nu}\Big]=0~.$ (43)
Eq. ( provides the dynamics of the fluid while Eq. ( describes its kinematics. In what follows, we will show that the analogous relations of Eq. ( and Eq. ( exist in gravitational dynamics on a generic null hypersurface, provided we pick out an adapted coordinate system w.r.t. $\mathcal{H}$. More importantly, those relations are obtained in a conventional manner i.e. taking different projections of conservation of suitably chosen gravitational energy-momentum tensor, similar as described in Eq. ( and Eq. ( for usual fluid.
We now discuss the notion of adapting a coordinate system with respect to our integrable null surface $\mathcal{H}$ as detailed out in [25]. As mentioned before, owing to the $3+1$ foliation of $\mathcal{H}$ by the family of spacelike surfaces $\Sigma_{t}$ allows us to coordinatize spacetime manifold in the vicinity of $\mathcal{H}$ by $(t,x^{1},x^{2},x^{3})$. Adapting the coordinates to the null surface means that we can fix the location of $\mathcal{H}$ by the choice $x^{1}=0$. Along the time evolution vector field $\bm{t}=\bm{\partial_{t}}$, the transverse/angular coordinates of $\mathcal{B}^{\rm(null)}$ are fixed. It can be shown [25], that on the null surface we have, $\bm{l}\overset{\mathcal{H}}{=}\bm{t}+\bm{V}$, where $\bm{V}$ is spatial vector field that is tangent to $\mathcal{B}^{\rm(null)}$. With respect to the adapted coordinate system, hence, we have $l^{\tilde{\mu}}\overset{\mathcal{H}}{=}(1,V^{A})$. The line element on the null surface in this adapted coordinate system can be expressed as [25, 26],
Now, it could be possible that for a specific choice of the adapted coordinates, we could have the constraint that $\partial_{t}q_{AB}=0$ i.e. the induced metric $q_{AB}$ is stationary or independent of the time evolution parameter $t$. Then the second fundamental form of the null surface $\mathcal{H}$ takes exactly the same structure as that of the stress tensor of a two-dimensional viscous fluid with velocity $V^{A}$. This part will be used to relate the gravitational equations with NS fluid equations Eq. ( and Eq. (.
For the case of Einstein gravity, upon analyzing the evolution equation for the Hajicek one-form Eq. ( in the adapted coordinate system w.r.t. $\mathcal{H}$ and using that $\pounds_{\bm{{l}}}\Omega_{a}\overset{\mathcal{H}}{=}\pounds_{\bm{t}}\Omega_{a}+\pounds_{\bm{V}}\Omega_{a}$, we have,
This is the well-known Damour-Navier-Stokes equation [23, 78]. In fact, it was shown by Padmanabhan [26], that analyzing the DNS equation Eq. ( with respect to a boosted local inertial frame, the extra term $\Omega_{b}~^{(2)}D_{a}V^{b}$ drops out and it takes the structure of a two dimensional viscous Navier-Stokes fluid, identical to Eq. ( for a specific choice of fluid parameters in terms of gravitational quantities. So following that idea, Eq. ( becomes (in the absence of external matter)